# Static Generation of Efficient OpenMP Offload Data Mappings

### v1

@@ Abstract
Increasing heterogeneity in HPC architectures and compiler advancements have led to OpenMP being frequently used to enable computations on heterogeneous devices. However, the efficient movement of data on heterogeneous computing platforms is crucial for achieving high utilization. The implicit OpenMP data-mapping rules often result in redundant data transfer, which can be a bottleneck for program performance. Programmers must explicitly map data between the host and connected accelerator devices to achieve efficient data movement. For this, OpenMP offers the target data and target update constructs. Ensuring efficient data transfer requires programmers to reason about complex data flow. This can be a laborious and error-prone process since the programmer must keep a mental model of data validity and lifetime spanning multiple data environments. Any automated analysis should maximize data reuse, minimize data transfer, and must consider control flow and context from function call sites, making the analysis interprocedural and context sensitive. In this paper, we present a static analysis tool, OMPDart (OpenMP Data Reduction Tool), for OpenMP programs that models data dependencies between host and device regions and applies source code transformations to achieve efficient data transfer. The analysis is based on a hybrid data structure that joins an Abstract Syntax Tree (AST) with a Control Flow Graph (CFG). Our evaluations on nine HPC benchmarks demonstrate that OMPDart is capable of generating effective data mapping constructs that substantially reduce data transfer between host and device. OMPDart helps reduce data transfers by $85\%$ and improves runtime performance by 1.6×1.6$\times$1.6 × over an expert defined implementation of LULESH $2.0$.
@@ I Introduction
In this work, we present static data-flow analysis techniques that include interprocedural and context-sensitive approaches to identify data dependencies between memory spaces. We introduce OMPDart, a static analysis tool for OpenMP programs designed to automate the data mapping process by modeling data dependencies between host and device regions and inserting data-mapping constructs in the source code to resolve them. Our techniques are based on a hybrid code representation that combines an Abstract Syntax Tree (AST) and a Control Flow Graph (CFG). Nine HPC benchmarks are used to evaluate the tool by comparing the performance of the tool-defined mappings against an expert hand-optimized implementation. We show that OMPDart can correctly generate data mappings that perform substantially better than default data mappings and on par with or in some cases better than the expert-defined mappings.
@@ IV-B Parsing the Abstract Syntax Tree
The AST and CFG are combined to form a hybrid AST-CFG representation in which each node of the CFG is linked with the corresponding AST representation. This joint data structure captures both the structural hierarchy of the code represented by the AST and the control flow relationships of the CFG. Figure 2 shows example code and the corresponding AST-CFG representation. The AST-CFG is constructed on a function level, similar to the process of constructing a Code Property Graph (CPG). CPGs were first introduced by Yamaguchi et al. [5] for the purposes of detecting vulnerabilities in C code.
@@ IV-C Interprocedural Analysis
Now that each procedure has been identified and has a CFG associated with it, an interprocedural analysis step is applied. This step analyzes how function calls affect the data of the caller. Specifically, we are interested in the passing of parameters and the use of common global variables across functions. The tool iteratively checks each function to determine if and in what manner any global variables or data at the address of pointer parameters were accessed. This includes whether data is read or written and if such accesses occur on the host or device. Once a determination is made about the use of pointer parameters or global variables, the model is augmented at each call site of the function with maximally pessimistic assumptions about the memory accesses of the callee. This process can be repeated several times up to the maximum call depth of any function. Each pass provides new information about the function to the model, but can be stopped early if no updates are made during a pass. This helps the model make more accurate assumptions about data accesses.
@@ IV-D Data-Flow Analysis
The analysis is performed as a forward traversal of the AST-CFG for each function. We trace the reads and writes to any variable referenced inside any offloaded region with the goal of identifying data dependencies between the host and device memory spaces. Anti (WAR) dependencies and output (WAW) dependencies do not necessitate any communication. Only true (RAW) dependencies need to be resolved by inserting one of the constructs described in Table II.
@@ IV-E Array Access Pattern Analysis
In [9], Guo et al. presented a compile-time algorithm that operates on the Clang AST and can successfully eliminate the transfer of unused array segments in C and C++ OpenMP programs. The algorithm makes determinations about what data transfer is truly necessary based on the analysis of for loops to determine which array bounds are accessed. This analysis may be impeded if the programmer’s use of a for statement lacks any of the initializing, conditional, or incrementing/decrementing statements, or if any of these statements are overly complex.
@@
OMPDart extends this bounds analysis to infer the access patterns in nested loops to assist the placement of OpenMP update directives. Consider the nested for loop from HeCBench’s OpenMP offloading implementation of the Rodinia backprop benchmark [11] shown in Listing 6. The code is executed on the host and reads from the array partial_sum, which happens to be invalid on the host upon entering the snippet. Using the logic presented in Section IV-D, the automated tool will correctly identify that somewhere proceeding line 5, it must insert the following OpenMP directive: target update from(partial_sum). The immediate choice would be to insert the statement between lines 4 and 5, however, this would cause the data transfer to occur every iteration of the inner loop. To avoid this redundant data transfer, the update directive should be placed before the outermost loop (preceding line 2). While both the options would produce technically correct code, in this example, placing the update directive in the inner loop would result in over $2$GB of data transfer compared to less than $5$MB when the directive is inserted before the for loops (with 65536 input elements). By correctly identifying the nested loop structure and access pattern of partial_sum, OMPDart can make a better decision on the placement of the update directive, which translates to a 14$\times$ speedup.
@@
Algorithm 1 can be used to find the outermost loop that impacts the indexing of a given array. $loops$ is a stack that contains references to for statements. The top of the stack corresponds to the for loop that includes array access $a$ in its body, with each subsequent element representing the enclosing loop. $locLim$ stores a statement which the directive must not proceed, typically being the end of the preceding target kernel’s scope. The procedure returns a pointer to the statement that the directive should directly proceed or follow depending on whether the direction of data movement is from or to the device, respectively. From there, this location is used to further decide where the best placement of the update to/from statement might be.

### v3

@@ Abstract
Increasing heterogeneity in HPC architectures and compiler advancements have led to OpenMP being frequently used to enable computations on heterogeneous devices. However, the efficient movement of data on heterogeneous computing platforms is crucial for achieving high utilization. Programmers must explicitly map data between the host and connected accelerator devices to achieve efficient data movement. Ensuring efficient data transfer requires programmers to reason about complex data flow. This can be a laborious and error-prone process since the programmer must keep a mental model of data validity and lifetime spanning multiple data environments. We present a static analysis tool, OMPDart (OpenMP Data Reduction Tool), for OpenMP programs that models data dependencies between host and device regions and applies source code transformations to achieve efficient data transfer. Our evaluations on nine HPC benchmarks demonstrate that OMPDart is capable of generating effective data mapping constructs that substantially reduce data transfer between host and device.
Index Terms:
High performance computing, Parallel programming, Data transfer, Static analysis, GPU, Performance optimization, Heterogeneous computing, OpenMP
@@ I Introduction
In this work, we present static data-flow analysis techniques that include interprocedural and context-sensitive approaches to identify data dependencies between memory spaces. We introduce OMPDart, a static analysis tool for OpenMP programs designed to automate the data mapping process by modeling data dependencies between host and device regions and inserting data-mapping constructs in the source code to resolve them. Our techniques are based on a hybrid code representation that combines an Abstract Syntax Tree (AST) and a Control Flow Graph (CFG). Nine HPC benchmarks are used to evaluate the tool by comparing the performance of the tool-defined mappings against an expert hand-optimized implementation. We show that OMPDart can correctly generate data mappings that perform substantially better than default data mappings and on par with or, in some cases, better than the expert-defined mappings.
@@ IV-B Parsing the Abstract Syntax Tree
The AST and CFG are combined to form a hybrid AST-CFG representation in which each node of the CFG is linked with the corresponding AST representation. To the best of our knowledge, this is a novel data structure, and its application in this context is unique. The AST-CFG is integral to our analysis since it maintains both the structural hierarchy of the code represented by the AST and the control flow relationships of the CFG. A traversal of the CFG edges models data flow, variable liveness, and data validity to determine data dependencies, while intermittent AST analysis assists in detecting array access patterns that are crucial for deciding when to transfer data. Figure 2 shows example code and the corresponding AST-CFG representation. The AST-CFG is constructed on a function level, similar to the process of constructing a Code Property Graph (CPG). CPGs were first introduced by Yamaguchi et al. [5] for the purposes of detecting vulnerabilities in C code.
@@ IV-C Interprocedural Analysis
Now that each procedure has been identified and has a CFG associated with it, an interprocedural analysis step is applied. This step analyzes how function calls affect the data of the caller. Specifically, we are interested in the passing of parameters and the use of common global variables across functions. The tool iteratively checks each function to determine if and in what manner any global variables or data at the address of pointer parameters were accessed. This includes whether data is read or written and if such accesses occur on the host or device. Once a determination is made about the use of pointer parameters or global variables, the model is augmented at each call site of the function with maximally pessimistic assumptions about the memory accesses of the callee. In this context, a maximally pessimistic approach means the analysis assumes the worst-case scenario for any unknown or uncertain information. This conservative approach helps to ensure that we don’t violate the correctness of a program, of course, with the potential downside that we may miss an optimal solution. This is particularly relevant in cases when we don’t have visibility into a function that may access memory in unknown ways. This process can be repeated several times up to the maximum call depth of any function. Each pass provides new information about the function to the model, but can be stopped early if no updates are made during a pass. This helps the model make more accurate assumptions about data accesses.
@@ IV-D Data-Flow Analysis
The analysis is performed as a forward traversal of the AST-CFG for each function. We trace the reads and writes to any variable referenced inside any offloaded region with the goal of identifying data dependencies between the host and device memory spaces. Anti (Write After Read, WAR) dependencies and output (Write After Write, WAW) dependencies do not necessitate any communication. Only true (Read After Write, RAW) dependencies need to be resolved by inserting one of the constructs described in Table II.
@@ IV-E Array Access Pattern Analysis
In [9], Guo et al. presented a compile-time algorithm that operates on the Clang AST and can successfully eliminate the transfer of unused array segments in C and C++ OpenMP programs. The algorithm makes determinations about what data transfer is truly necessary based on the analysis of for loops to determine which array bounds are accessed. This analysis may be impeded if the programmer’s use of a for statement lacks any of the initializing, conditional, or incrementing/decrementing statements, or if any of these statements are overly complex. Furthermore, the algorithm presented in [9] has no mechanism to work with nested loops and uses the lifetime of a for loop as the lifetime to map the data. In this section, we describe how we modify their approach for the purposes of arrays with multiple dimensions and nested loops, and how we determine efficient placement of update directives in scenarios involving nested loops.
@@
OMPDart extends this bounds analysis to infer the access patterns in nested loops to assist the placement of OpenMP update directives. Consider the nested for loop from HeCBench’s OpenMP offloading implementation of the Rodinia backprop benchmark [11] shown in Listing 6. The code is executed on the host and reads from the array partial_sum, which happens to be invalid on the host upon entering the snippet. Using the logic presented in Section IV-D, the automated tool will correctly identify that somewhere preceding line 5, it must insert the following OpenMP directive: target update from(partial_sum). The immediate choice would be to insert the statement between lines 4 and 5, however, this would cause the data transfer to occur every iteration of the inner loop. To avoid this redundant data transfer, the update directive should be placed before the outermost loop (preceding line 2). While both the options would produce technically correct code, in this example, placing the update directive in the inner loop would result in over $2$GB of data transfer compared to less than $5$MB when the directive is inserted before the for loops (with 65536 input elements). By correctly identifying the nested loop structure and access pattern of partial_sum, OMPDart can make a better decision on the placement of the update directive, which translates to a 14$\times$ speedup.
@@
Algorithm 1 can be used to find the outermost loop that impacts the indexing of a given array. $loops$ is a stack that contains references to for statements. The top of the stack corresponds to the for loop that includes array access $a$ in its body, with each subsequent element representing the enclosing loop. $locLim$ stores a statement which the directive must not precede, typically being the end of the preceding target kernel’s scope. The procedure returns a pointer to the statement that the directive should directly precede or follow depending on whether the direction of data movement is from or to the device, respectively. From there, this location is used to further decide where the best placement of the update to/from statement might be.
@@ VI Experiments
The tool’s execution time is negligible. The average execution time of OMPDart on the benchmarks studied was $0.29$s. lulesh, with 15 kernels, was the biggest benchmark evaluated and had the greatest overhead of $1.35$s. Table V lists the tool’s execution time to analyze and transform each benchmark.
TABLE V: OMPDart Overhead
Benchmark Tool Execution Time
accuracy 0.34s
ace 0.18s
backprop 0.34s
bfs 0.14s
clenergy 0.16s
hotspot 0.05s
lulesh 1.35s
nw 0.14s
xsbench 0.14s
@@ VII Discussion
At the time of our evaluation, AMD GPUs were not readily available to us, hence they were not tested. However, we believe a similar reduction in data transfers should be seen on AMD GPUs as well.
@@ IX Conclusion
Acknowledgements
This research was supported by the National Science Foundation under Grant number 2211982. We would also like to thank the Research IT team [^1] https://researchit.las.iastate.edu/ of Iowa State University for their support and for providing access to the HPC Cluster, which were essential for conducting the experiments in this project.